Prove that if an integer n is not divisible by 3, then n^2=3k+1 for some integer k.

Note: “not divisible by 3” means either “n=3m+1 for some integer m” or “n=3m+2 for some integer m”.
Answer #1

Assume that n is at not divisible by 3 > n = 3m+ or 3m +2 for some mez Casei N= 3m+1 = n? (3m+1)?= gm?+ 6M+/ = 3(3m?42m) +1 T